The information below applies to the day of investment in a mutual fund:


Amount
Invested

Total Fund
Assets

Total Fund
Liabilities

Total Shares
Outstanding

$3700 $761 million $48 million 75 million

Find the net asset value and the number of shares purchased.

Answer by Aldorozos(172)

Here is to the best of my knowledge
We subtract the liability from the total assets to calculate net asset value
761-48 = 713 mil.
We divide the net asset value to the total number of shares to see how much is the value of each share
713/75 = 9.507
Now that we are investing $3700, we have to see how many shares we can purchase
$3700/9.507 = 389 shares